OpenStudy (anonymous):

The given measurements may or may not determine a triangle. If not, then state that no triangle is formed. If a triangle is formed, then use the Law of Sines to solve the triangle, if it is possible, or state that the Law of Sines cannot be used. B = 137°, c = 9, b = 14

OpenStudy (anonymous):

what do u need us to do ? evaluate?

OpenStudy (anonymous):

These are the answer options. They want you to evaluate the measurements given above i think. -No triangle is formed. -C = 26°, A = 17°, a ≈ 6 -The triangle cannot be solved with the Law of Sines. -C = 17°, A = 26°, a ≈ 6

OpenStudy (anonymous):

ok i got u

OpenStudy (anonymous):

so? @JULIANCHICO

OpenStudy (raden):

b sinC = c sinB sinC = (c sinB)/b sinC = (9 * sin137)/14 = 0.438 <C = sin^-1(0.438) = 26 deg

OpenStudy (raden):

<A + <B + <c = 180 <A = 180 - (<B+<C) <A = 180 - (137+26) <A = 180 - 163 <A = 17 deg

OpenStudy (raden):

a sinB = b sinA a = b sinA/sinB a = 14 sin17/sin137 a = 14 * 0.292/0.682 = 5.99... (take nearest = 6)